OpenStudy (anonymous):

A four sided, balanced die has the values 1, 2, 3, and 4. There are three successive throws. Find the probability of throwing the same value on exactly two throws

ganeshie8 (ganeshie8):

Total possible permutations = \(4*4*4 = 4^3\) Next observe that you can fix same number on two dice in \(4\) ways, After that, the third die can be any of remaining \(3\) numbers So there are \(4\times 3 = 12\) ways for this to happen and accounting for permutations we have \(12\times 3 = 36\) total favorable outcomes.

ganeshie8 (ganeshie8):

Take the ratio of favorable and total for the probability : \[\dfrac{36}{4^3}\]

ganeshie8 (ganeshie8):

You could also work it by considering a 3 bit string. you can choose first digit to be anything so probability is 1 after that, you may choose the second digit to be same as first digit : 1/4 Finally the third digit must be different : 3/4 So the probability for having first two same digits and last digit different would be \[1\cdot \dfrac{1}{4}\cdot \dfrac{3}{4}\] Multiply that by \(3\) to account for order

OpenStudy (mathmate):

Another way is to subtract from 1 the probabilities having all different numbers (4*3*2/64) and all identical (4/64). 1-(4*3*2)/64-4/64=36/64

OpenStudy (anonymous):

Ahhh okay, I understand it now! It was driving me crazy lol. Thanks for the help!